Divaseventeen:

If a rectangular lot measures 160 ft. by 110 ft. and 15 foot set backs are required from each lot line, how much area is left? If 65 cars require 32,500 square feet, how many cars will the small rectangle hold? There will be square feet in the smaller lot. (If applicable, use a thousands comma.) The number of cars is . (Enter a whole number.)

sarah123:

tbh idk

SandMandGirl:

(160 - 2 x 15) (110 - 2 x15) = 130 x 80 = 10400 There will be 10400 square feet in the smaller lot. \[\frac{ 32500 }{ 65 }\] = 500 square feet per car \[\frac{ 10400 }{ 500 }\] = 20.8 ~ 20 The number of cars is 20